OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 4x – 5y = 24? Choose all answers that are correct. A. (−9, −12) B. (1, −4) C. (4, 8) D. (6, 0) @ganeshie8 @goformit100 @jackoo @ozzy22 @1DEA @tilt @Jesstho.-.

OpenStudy (anonymous):

A,B,D

ganeshie8 (ganeshie8):

thats right, could you also explain how you figured that out ?

OpenStudy (anonymous):

can you guys help me with this one Which ordered pairs in the form (x, y) are solutions to the equation (x + 3)y = 14? Choose all answers that are correct. A. (−5, −7) B. (−1, 7) C. (10, −2) D. (11, 1) @ganeshie8

OpenStudy (anonymous):

Plug them in.

OpenStudy (anonymous):

4x – 5y = 24 4(-9)-5(-12)=24 -36+60=24 24=24 <---good 4(1)-5(-4)=24 4+20=24 24=24<---good 4(4)-5(8)=24 16-40=24 -24=24<---no 4(6)-5(0)=24 24=24<---good

OpenStudy (anonymous):

(x + 3)y = 14 (-5+3)7 = 14 -35+21 = 14 -14 = 14<---no (-1+3)7=14 -7+23=14 16=14<---no (10+3)-2=14 -20-6=14 -26=14<---no (11+3)1=14 11+3=14 14=14<----yes So, D